Now that they have accumulated a deposit of $100,000, Joe and Jenny wish to use the deposit and take out a housing loan to purchase a home. The home costs $900,000. The loan is to be repaid in equal monthly instalments over a term of 20 years. Jenny recalls that the interest rate quoted by the bank is an annual nominal rate of 6.0%pa compounded monthly. After 5 years (60th repayment just made), the bank announces the interest rate will change to an annual nominal rate of 7.2%pa compounded monthly.

(i) How much is the original monthly repayment?

(ii) How much is the monthly repayment after the interest rate rise, assuming the term is unchanged?

(ii) Provide Joe and Jenny with a repayment schedule after the interest rate rise using excel.
Answer #1

Solution i) Calculation of Original Monthly repayment installment

Following are the steps to be followed on Microsoft Excel to calculate the Monthly Installment:

Step 1: Click on "FORMULAS" tab at the top of Microsoft Excel
Step 2: Select the option "Financial"
Step 3: Under "Financial" select the option "PMT"
Step 4: Insert Rate = 0.06/12 Nper = 20*12 PV = -800000

PMT = $5731.45

Therefore, the original monthly installment is $5731.45.

Solution ii) Calculation of Monthly repayment installment after the change in Interest rate to 7.2% per year

After 60th Installment, the outstanding loan = $679196.78

Following are the steps to be followed on Microsoft Excel to calculate the Interest payment:

Step 1: Click on "FORMULAS" tab at the top of Microsoft Excel
Step 2: Select the option "Financial"
Step 3: Under "Financial" select the option "PMT"
Step 4: Insert Rate = 0.072/12    Nper = 15*12 PV = -679196.78

PMT = $6181.01

Therefore, the new monthly installment after 60th Month is $6181.01.

Solution iii) Loan Repayment Schedule

Following are the steps to be followed on Microsoft Excel to calculate the Interest payment:

Step 1: Click on "FORMULAS" tab at the top of Microsoft Excel
Step 2: Select the option "Financial"
Step 3: Under "Financial" select the option "IPMT"
Step 4: Insert Rate = 0.072/12    Per = 1 Nper = 15*12 PV = -679196.78

IPMT = $4075.18

Principal payment = Total payment – Interest payment

Ending balance = Beginning balance – Principal payment

Keeping the Rate, Nper and PV as constant drag the formula till 180th Month, you will derive the schedule

The Entire repayment schedule cannot be pasted. Hence, I am providing the repayment schedule for next 5 years.

Month

Beginning Balance

Total Payment

Interest Payment

Principal Payment

Ending Balance

61

6,79,196.79

6,181.01

4,075.18

2,105.83

6,77,090.96

62

6,77,090.96

6,181.01

4,062.55

2,118.46

6,74,972.50

63

6,74,972.50

6,181.01

4,049.83

2,131.17

6,72,841.32

64

6,72,841.32

6,181.01

4,037.05

2,143.96

6,70,697.36

65

6,70,697.36

6,181.01

4,024.18

2,156.82

6,68,540.54

66

6,68,540.54

6,181.01

4,011.24

2,169.76

6,66,370.78

67

6,66,370.78

6,181.01

3,998.22

2,182.78

6,64,187.99

68

6,64,187.99

6,181.01

3,985.13

2,195.88

6,61,992.11

69

6,61,992.11

6,181.01

3,971.95

2,209.06

6,59,783.06

70

6,59,783.06

6,181.01

3,958.70

2,222.31

6,57,560.75

71

6,57,560.75

6,181.01

3,945.36

2,235.64

6,55,325.10

72

6,55,325.10

6,181.01

3,931.95

2,249.06

6,53,076.04

73

6,53,076.04

6,181.01

3,918.46

2,262.55

6,50,813.49

74

6,50,813.49

6,181.01

3,904.88

2,276.13

6,48,537.37

75

6,48,537.37

6,181.01

3,891.22

2,289.78

6,46,247.58

76

6,46,247.58

6,181.01

3,877.49

2,303.52

6,43,944.06

77

6,43,944.06

6,181.01

3,863.66

2,317.34

6,41,626.71

78

6,41,626.71

6,181.01

3,849.76

2,331.25

6,39,295.47

79

6,39,295.47

6,181.01

3,835.77

2,345.24

6,36,950.23

80

6,36,950.23

6,181.01

3,821.70

2,359.31

6,34,590.92

81

6,34,590.92

6,181.01

3,807.55

2,373.46

6,32,217.46

82

6,32,217.46

6,181.01

3,793.30

2,387.70

6,29,829.76

83

6,29,829.76

6,181.01

3,778.98

2,402.03

6,27,427.73

84

6,27,427.73

6,181.01

3,764.57

2,416.44

6,25,011.29

85

6,25,011.29

6,181.01

3,750.07

2,430.94

6,22,580.35

86

6,22,580.35

6,181.01

3,735.48

2,445.53

6,20,134.82

87

6,20,134.82

6,181.01

3,720.81

2,460.20

6,17,674.62

88

6,17,674.62

6,181.01

3,706.05

2,474.96

6,15,199.66

89

6,15,199.66

6,181.01

3,691.20

2,489.81

6,12,709.85

90

6,12,709.85

6,181.01

3,676.26

2,504.75

6,10,205.10

91

6,10,205.10

6,181.01

3,661.23

2,519.78

6,07,685.32

92

6,07,685.32

6,181.01

3,646.11

2,534.90

6,05,150.43

93

6,05,150.43

6,181.01

3,630.90

2,550.11

6,02,600.32

94

6,02,600.32

6,181.01

3,615.60

2,565.41

6,00,034.92

95

6,00,034.92

6,181.01

3,600.21

2,580.80

5,97,454.12

96

5,97,454.12

6,181.01

3,584.72

2,596.28

5,94,857.83

97

5,94,857.83

6,181.01

3,569.15

2,611.86

5,92,245.97

98

5,92,245.97

6,181.01

3,553.48

2,627.53

5,89,618.44

99

5,89,618.44

6,181.01

3,537.71

2,643.30

5,86,975.14

100

5,86,975.14

6,181.01

3,521.85

2,659.16

5,84,315.98

101

5,84,315.98

6,181.01

3,505.90

2,675.11

5,81,640.87

102

5,81,640.87

6,181.01

3,489.85

2,691.16

5,78,949.71

103

5,78,949.71

6,181.01

3,473.70

2,707.31

5,76,242.40

104

5,76,242.40

6,181.01

3,457.45

2,723.55

5,73,518.85

105

5,73,518.85

6,181.01

3,441.11

2,739.90

5,70,778.95

106

5,70,778.95

6,181.01

3,424.67

2,756.33

5,68,022.62

107

5,68,022.62

6,181.01

3,408.14

2,772.87

5,65,249.74

108

5,65,249.74

6,181.01

3,391.50

2,789.51

5,62,460.23

109

5,62,460.23

6,181.01

3,374.76

2,806.25

5,59,653.99

110

5,59,653.99

6,181.01

3,357.92

2,823.08

5,56,830.90

111

5,56,830.90

6,181.01

3,340.99

2,840.02

5,53,990.88

112

5,53,990.88

6,181.01

3,323.95

2,857.06

5,51,133.82

113

5,51,133.82

6,181.01

3,306.80

2,874.21

5,48,259.61

114

5,48,259.61

6,181.01

3,289.56

2,891.45

5,45,368.16

115

5,45,368.16

6,181.01

3,272.21

2,908.80

5,42,459.36

116

5,42,459.36

6,181.01

3,254.76

2,926.25

5,39,533.11

117

5,39,533.11

6,181.01

3,237.20

2,943.81

5,36,589.30

118

5,36,589.30

6,181.01

3,219.54

2,961.47

5,33,627.83

119

5,33,627.83

6,181.01

3,201.77

2,979.24

5,30,648.59

120

5,30,648.59

6,181.01

3,183.89

2,997.12

5,27,651.47